OpenStudy (anonymous):

Need help Mark is in a deep hole looking for treasure. He is standing 8 feet below the surface. He throws an old coin he found with an initial upward velocity of 22 ft/sec. How long until it lands outside the hole, having gone up and come back down? Use the formula h=-16t^2+22t-8, where h is the height of the coin in feet (relative to the surface) and t is the time in seconds since Mark threw it. Ignore air resistance and round your answer to the nearest tenth. A. 0.3 seconds B. 1.7 seconds C. 2.1 seconds D. It will not make it outside the hole.

OpenStudy (misssunshinexxoxo):

What do you believe the answer is?

OpenStudy (anonymous):

IT has to be D i've been stuck in this question for a while

OpenStudy (misssunshinexxoxo):

Sounds like physics I'll help you

OpenStudy (misssunshinexxoxo):

h = -16t^2+22t-8 h' = -32t + 22 = 0 t = 11/16 h = -16 (11/16)^2 +22(11/16) - 8 h = -0.4375 it will not make it out of the hole.

OpenStudy (misssunshinexxoxo):

Yes you are correct! :) Way to go

OpenStudy (anonymous):

Thanks